OpenStudy (anonymous):

Simplify each expression. Rationalize denominators. Show all work, step-by-step. sqrt w^4/x^3

OpenStudy (anonymous):

Is the bottom a sqrt too?

OpenStudy (anonymous):

If so, multiply top and bottom by sqrt x^3

OpenStudy (anonymous):

this is what it looks like

OpenStudy (anonymous):

Right well that is the same as \[\sqrt{w^4}/\sqrt{x^3}\] So just multiply the top and the bottom by \[\sqrt{x^3}\] giving you x^3 on the bottom and \[\sqrt{w^4}\sqrt{x}\] on the top

OpenStudy (anonymous):

thats it or more to it

OpenStudy (anonymous):

help @tcarroll010

OpenStudy (anonymous):

Multiply the fraction by root x^3 over root x^3

OpenStudy (anonymous):

\[\frac{ \sqrt{w^4}\sqrt{x^3} }{ x^3 }\]

OpenStudy (anonymous):

And sqrt of w^4 is just w^2 so \[\frac{ w^2\sqrt{x^3} }{ x^3 }\]

OpenStudy (anonymous):

That's it.

OpenStudy (anonymous):

thank u!!